(0.9.03-0.1244 log (60.5-31)/(0.9255-0.0979 log (60.5-31))-0.0839


1
Expert's answer
2021-08-16T08:39:00-0400

Solve the equation:

(0.9.03-0.1244 log (60.5-31)/(0.9255-0.0979 log (60.5-31))-0.0839.

We see that there is no closing parenthesis in the first part:

(0.9.03-0.1244 log (60.5-31)[possibly missing parenthesis]/(0.9255-0.0979 log (60.5-31))-0.0839.

So, to solve this equation, we need to omit the first parenthesis.

Also, we need to omit one of the periods in the very first term 0.9.03. It seems logical to use 0.903 instead:



"0.903-\frac{0.1244 \log (60.5-31)}{0.9255-0.0979 \log (60.5-31)}-0.0839=\\\space\\=0.5852."